# Steer Agent Ethical Behavior with Model Editing

### v1

@@ Abstract
Agents based on Large Language Models (LLMs) have demonstrated strong capabilities across a wide range of tasks. However, deploying LLM-based agents in high-stakes domains comes with significant safety and ethical risks. Unethical behavior by these agents can directly result in serious real-world consequences, including physical harm and financial loss. To efficiently steer the ethical behavior of agents, we frame agent behavior steering as a model editing task, which we term Behavior Editing. Model editing is an emerging area of research that enables precise and efficient modifications to LLMs while preserving their overall capabilities. To systematically study and evaluate this approach, we introduce BehaviorBench, a multi-tier benchmark grounded in psychological moral theories. This benchmark supports both the evaluation and editing of agent behaviors across a variety of scenarios, with each tier introducing more complex and ambiguous scenarios. We first demonstrate that Behavior Editing can dynamically steer agents toward the target behavior within specific scenarios. Moreover, Behavior Editing enables not only scenario-specific local adjustments but also more extensive shifts in an agent’s global moral alignment. We demonstrate that Behavior Editing can be used to promote ethical and benevolent behavior or, conversely, to induce harmful or malicious behavior. Through comprehensive evaluations on agents based on frontier LLMs, BehaviorBench shows the effectiveness of Behavior Editing across different models and scenarios. Our findings offer key insights into a new paradigm for steering agent behavior, highlighting both the promise and perils of Behavior Editing. Our code, data, and results are available at https://github.com/baixianghuang/behavior-edit
@@ 1 Introduction
Agents based on Large Language Models (LLMs) have become increasingly capable of performing a wide range of complex tasks (Guo et al., 2024). As these agents are increasingly deployed in high-stakes domains such as healthcare, finance, and education, they exert a direct and consequential influence on real-world decisions and outcomes (Xi et al., 2025). However, this progress is accompanied by serious concerns regarding the safety and ethical reliability of agent systems (Bengio et al., 2025). Unethical behavior by agents can lead to serious real-world consequences, including physical harm, financial loss, and erosion of public trust (Gan et al., 2024). Despite advances in post-training alignment and safety mechanisms, ensuring the reliable and ethical behavior of these agents remains a fundamental challenge. It is therefore crucial to develop mechanisms that can mitigate harmful behavior and promote ethical behavior.
Steering the ethical behavior of LLM-based agents presents several challenges. The first is the difficulty of measuring and quantifying ethical behavior in a systematic and principled way (Hendrycks et al., 2020, 2021). Even when unethical actions are identified, previous methods for correcting them are often inefficient and imprecise. Existing safeguards, such as full-parameter fine-tuning or hard-coded rules, often fall short in dynamic or context-dependent situations where ethical reasoning is nuanced and evolving (Bai et al., 2022; Sharma et al., 2025). As for moral alignment techniques, such as reinforcement learning from human feedback (RLHF) (Ouyang et al., 2022), they typically occur at the post-training stage and focus on broad alignment with human values. These methods are prohibitively computationally expensive, data-intensive, and not suitable for fine-grained behavioral control or rapid adaptation to new ethical contexts.
Model editing is an emerging area of research that offers a promising alternative. It allows efficient and targeted modifications of language models while minimizing disruptions to their overall knowledge and capabilities (Meng et al., 2022; Wang et al., 2024b; Zhang et al., 2024). While existing work on model editing has focused primarily on updating factual knowledge, its demonstrated effectiveness in making precise and accurate changes to factual knowledge motivates us to extend this idea to ethical behavior steering and introduce the concept of Behavior Editing, which enables directional steering of agent behavior through editing either the agent’s actions or its underlying moral judgments. Behavior Editing not only allows for scenario-specific behavioral adjustments but also enables more extensive shifts in the agent’s global moral alignment. Crucially, this capacity to steer behavior operates in both directions as shown in Figure 1: it can be used to promote benevolent behavior or to induce harmful behavior. In this sense, Behavior Editing is a double-edged sword, simultaneously enabling beneficial interventions and posing significant safety risks.
To systematically investigate this emerging paradigm, we introduce BehaviorBench, a multi-tier benchmark designed to evaluate the effectiveness of behavior editing techniques. Grounded in psychological theories including Moral Foundations Theory (Graham et al., 2013), Stages of Moral Development (Kohlberg, 1971), Normative Ethics (Kagan, 2018), and Rest’s Four Component Model (Narvaez & Rest, 1995), BehaviorBench includes a set of representative model editing approaches, an evaluation framework, and a curated collection of ethical scenarios and dilemmas as summarized in Table 1. Through comprehensive experiments across agents based on both proprietary and open-weight LLMs, we demonstrate that Behavior Editing enables reliable steering of agent behavior across diverse scenarios. Our findings provide new insights into the promise and perils of Behavior Editing, highlighting its potential to enable safer, more ethical agent systems while also revealing the serious risks associated with its misuse.
@@
Tier 1: Moral Sensitivity Detecting moral relevance in everyday situations, grounded in moral sensitivity theory, pre-conventional reasoning, and social norms. Social Chemistry 101
@@
Table 1: Three-tier structure of the BehaviorBench ethical behavior evaluation benchmark. As tiers progress from Moral Sensitivity to Moral Judgment and Moral Agency, scenarios become increasingly complex and cognitively demanding, reflecting a progression through Rest’s moral development model (moral sensitivity, moral judgment, motivation and character) (Narvaez & Rest, 1995), Kohlberg’s Stages of Moral Development (pre-conventional, conventional, post-conventional stage) (Kohlberg, 1971), and Normative Ethics (Kagan, 2018).
@@ 2.1 Model Editing
Model editing, also known as knowledge editing (Wang et al., 2024b), has emerged as a critical area of research for modifying LLMs without the need for large datasets or costly retraining. These methods enable precise and efficient updates to specific knowledge while preserving overall model capabilities. Model editing approaches can be broadly categorized into two groups: parameter-modifying and parameter-preserving methods. Parameter-modifying methods alter the model’s internal weights to encode new knowledge or behavior. This includes Locate-then-edit techniques such as ROME (Meng et al., 2022) and MEMIT (Meng et al., 2023), which identify relevant knowledge within the model before applying targeted modifications, and Constrained Fine-Tuning (Zhu et al., 2020; Meng et al., 2022; Zhang et al., 2024), which selectively fine-tunes specific layers of the model while minimizing unintended changes. In contrast, parameter-preserving methods such as In-Context Editing (ICE) (Zheng et al., 2023) avoid any modification to model parameters by embedding the desired information directly into the input context at inference time, enabling flexible and temporary behavior shifts without altering the underlying model.
These editing techniques have demonstrated effectiveness in updating factual knowledge and reducing hallucinations (Zhang et al., 2024; Huang et al., 2025). However, they also raise safety concerns, particularly the risk of injecting harmful or misleading content (Wang et al., 2024a; Chen et al., 2024). As such, model editing presents both opportunities and challenges in the pursuit of more reliable and controllable LLMs. Compared to prior work, we demonstrate that model editing can be an effective and precise method for steering LLM-based agents toward specific actions. Furthermore, we find that behavior editing has a substantial impact on the model’s global moral alignment.
@@ 2.2 Ethical Behavior of LLM-based Agents
Datasets for machine ethics and morality include Social Chemistry (Forbes et al., 2020), ETHICS (Hendrycks et al., 2020), and Jiminy Cricket (Hendrycks et al., 2021). Our proposed BehaviorBench framework organizes and improves existing datasets to evaluate agents across three tiers of moral competence grounded in psychological and philosophical theories: Moral Sensitivity (detecting ethical issues), Moral Judgment (making and justifying moral decisions), and Moral Agency (acting and reasoning in ambiguous dilemmas). Unlike existing benchmarks that primarily measure harm avoidance, BehaviorBench comprehensively assesses ethical reasoning capabilities through the lens of established normative frameworks (Kagan, 2018) including deontology, utilitarianism, virtue ethics, justice theory, and commonsense morality, providing a more nuanced understanding of how editing techniques can steer agent behavior toward various ethical perspectives.
@@
Datasets relevant to machine ethics and moral reasoning include Social Chemistry (Forbes et al., 2020), MoralChoice (Scherrer et al., 2023), ETHICS (Hendrycks et al., 2020), and Jiminy Cricket (Hendrycks et al., 2021). Building on these foundations, our proposed BehaviorBench systematically organizes and enhances existing datasets to evaluate agents across three tiers of moral competence, grounded in psychological and philosophical theory: Moral Sensitivity (recognition of ethical issues), Moral Judgment (reasoned decision-making and justification), and Moral Agency (deliberation and action in ethically ambiguous contexts). In comparison to prior benchmarks that primarily emphasize harm avoidance, BehaviorBench offers a more comprehensive assessment of ethical reasoning. It also incorporates a range of normative ethical theories, including deontology, utilitarianism, virtue ethics, theories of justice, and commonsense morality, following the principles of Normative Ethics Kagan (2018). This multidimensional approach facilitates a more nuanced understanding of how editing techniques can steer agent behavior toward specific ethical orientations.
@@
Various approaches have been proposed to instill ethical constraints and guide the behavior of LLM-based agents. Recent work has focused primarily on alignment techniques. Techniques such as reinforcement learning from human feedback (RLHF) (Ouyang et al., 2022) enable LLMs to align closely with human ethical intuitions by learning from explicit human judgments. Constitutional AI (Sharma et al., 2025; Bai et al., 2022) extends this by allowing models to critique their own outputs against a set of principles. Other methodologies leverage rule-based ethical frameworks or explicit fine-tuning to embed ethical principles directly into model parameters (Choi et al., 2024).
@@
Unlike traditional alignment methods that tune the overall behavior of the model, model editing offers more precise intervention by targeting specific knowledge or response patterns within the parameters of the model while preserving other capabilities (Meng et al., 2022; Zhang et al., 2024). Where RLHF and similar approaches require extensive datasets, computationally expensive retraining, and human involvement, behavior editing can implement targeted behavioral changes with significantly lower computational overhead. This surgical precision makes model editing particularly well-suited for steering ethical behavior in complex scenarios where general alignment techniques might over-constrain agent behavior or fail to address nuanced ethical distinctions.
@@ 3.1 Problem Formulation
The goal of Behavior Editing is to precisely and efficiently steer the behavior of LLM-based agents while preserving their general capabilities. Behavior Editing has two primary directions: Benevolent Behavior Editing, which enhances positive behaviors by steering agents toward more friendly, helpful, and altruistic responses, and Malicious Behavior Editing, which deliberately introduces harmful behaviors to manipulate agents into behaving selfishly and hostilely, thereby compromising their moral and safety alignment.
@@
Behavior Editing operates on a structure analogous to a knowledge tuple $(s,r,o)$, where traditionally $s$, $r$, and $o$ denote the subject, relation, and object, respectively. However, in the context of Behavior Editing, the interpretations of $s$ and $r$ vary depending on the specific editing setting. We distinguish between two primary categories: Behavior-as-target editing and Judgment-as-target editing, both of which can be represented using the same tuple notation for consistency. In the Behavior-as-target setting, the goal is to modify a behavior exhibited in a given moral scenario. This is formalized as transforming an original tuple $(s,r,o)$, where $s$ is a hypothetical moral scenario, $r$ is the relation to behavior, and $o$ is the behavior under that scenario, into a new tuple $(s,r,o^{*})$ that reflects the edited behavior. Here, the scenario remains constant while the behavior changes. In contrast, Judgment-as-target editing focuses on altering the moral judgment associated with a given behavior. This is represented as transforming the tuple $(s,r,o)$, where $s$ denotes a behavior, $r$ is the relation to moral evaluation, and $o$ is the original moral judgment, into $(s,r,o^{*})$, where $o^{*}$ is the updated judgment. In both cases, an editing operation can be compactly expressed as $e=(s,r,o,o^{*})$, capturing the transformation from the original to the modified output.
@@
To analyze and modify an agent’s behavior in a given scenario, the scenario must first be converted into a natural language question $x$, to which the agent responds with an answer $y$. This input-output pair is associated with a behavior tuple $(s,r,o)$. The input space corresponding to an edit is denoted as $\mathcal{X}_{e}=I(s,r)$, where $I$ maps the scenario and relation to a set of relevant inputs. The original output space is defined as $\mathcal{Y}_{e}=O(s,r,o)$, and the desired or target output space after editing is represented as $\mathcal{Y}_{e}^{*}=O^{*}(s,r,o^{*})$. For a single edit $e$ with input space $\mathcal{X}_{e}$, the objective of Behavior Editing is to transform the original outputs $\mathcal{Y}_{e}$ into the target outputs $\mathcal{Y}_{e}^{*}$. When considering a set of edits =$\mathcal{E}=\{e_{1},e_{2},\ldots\}$, the combined input space is $\mathcal{X}_{\mathcal{E}}=\bigcup_{e\in\mathcal{E}}\mathcal{X}_{e}$, and the corresponding original and target output spaces are $\mathcal{Y}_{\mathcal{E}}=\bigcup_{e\in\mathcal{E}}\mathcal{Y}_{e}$ and $\mathcal{Y}_{\mathcal{E}}^{*}=\bigcup_{e\in\mathcal{E}}\mathcal{Y}_{e}^{*}$, respectively.
@@
The overarching goal of Behavior Editing is to modify an LLM-based agent, initially represented as a function $f:\mathcal{X}\rightarrow\mathcal{Y}$, into a new function $f^{*}:\mathcal{X}\rightarrow\mathcal{Y}^{*}$, such that the edited model generates the target behavior for inputs in $\mathcal{X}_{\mathcal{E}}$ while preserving its behavior on all other inputs. The optimization aims to minimize the discrepancy between the edited output $f^{*}(x)$ and the desired behavior $y^{*}$, as measured by a loss function $\mathcal{L}$. At the same time, the editing must maintain consistency on all inputs outside the editing set, ensuring that $f^{*}(x)=f(x)$ for all $x\in\mathcal{X}\setminus\mathcal{X}_{\mathcal{E}}$. This leads to the following constrained optimization objective:
@@
[…]

### v2

# Steering Agent Ethical Behavior Toward Beneficence or Harm

@@ Abstract
Agents based on Large Language Models (LLMs) have demonstrated strong capabilities across a wide range of tasks. However, deploying LLM-based agents in high-stakes domains comes with significant safety and ethical risks. Unethical behavior by these agents can directly result in serious real-world consequences, including physical harm and financial loss. To efficiently steer the ethical behavior of agents, we frame agent behavior steering as a model editing task, which we term Behavior Editing. Model editing is an emerging area of research that enables precise and efficient modifications to LLMs while preserving their overall capabilities. To systematically study and evaluate this approach, we introduce BehaviorBench, a multi-tier benchmark grounded in psychological moral theories. This benchmark supports both the evaluation and editing of agent behaviors across a variety of scenarios, with each tier introducing more complex and ambiguous scenarios. We first demonstrate that Behavior Editing can dynamically steer agents toward the target behavior within specific scenarios. Moreover, Behavior Editing enables not only scenario-specific local adjustments but also more extensive shifts in an agent’s global moral alignment. We demonstrate that Behavior Editing can be used to promote ethical and benevolent behavior or, conversely, to induce harmful or malicious behavior. Through extensive evaluations of agents built on frontier LLMs, BehaviorBench validates the effectiveness of behavior editing across a wide range of models and scenarios. Our findings offer key insights into a new paradigm for steering agent behavior, highlighting both the promise and perils of Behavior Editing.
@@
Websitehttps://model-editing.github.io
Codehttps://github.com/baixianghuang/behavior-edit
@@ 1 Introduction
Agents based on Large Language Models (LLMs) have become increasingly capable of performing a wide range of complex tasks (Guo et al. 2024). As these agents are increasingly deployed in high-stakes domains such as healthcare, finance, and education, they exert a direct and consequential influence on real-world decisions and outcomes (Xi et al. 2025). However, this progress is accompanied by serious concerns regarding the safety and ethical reliability of agent systems (Bengio et al. 2025). Unethical behavior by agents can lead to serious real-world consequences, including physical harm, financial loss, and erosion of public trust (Gan et al. 2024). Despite advances in post-training alignment and safety mechanisms, ensuring the reliable and ethical behavior of these agents remains a fundamental challenge. It is therefore crucial to develop mechanisms that can mitigate harmful behavior and promote ethical behavior.
Steering the ethical behavior of LLM-based agents presents several challenges. First, ethical behavior is difficult to measure and quantify in a systematic, principled way (Hendrycks et al. 2020). Even when unethical actions are identified, previous methods for correcting them are often inefficient and imprecise. Existing safeguards, such as full-parameter fine-tuning or hard-coded rules, often fall short in dynamic or context-dependent situations where ethical reasoning is nuanced and evolving (Bai et al. 2022; Sharma et al. 2025). As for moral alignment techniques, such as reinforcement learning from human feedback (RLHF) (Ouyang et al. 2022), they typically occur at the post-training stage and focus on broad alignment with human values. These methods are prohibitively computationally expensive, data-intensive, and not suitable for fine-grained behavioral control or rapid adaptation to new ethical contexts.
@@
Tier 1: Moral Sensitivity Detecting moral relevance, grounded in moral sensitivity theory, pre-conventional reasoning, and social norms. Social Chemistry 101
@@
Table 1: Three-tier structure of the BehaviorBench ethical behavior evaluation benchmark. As tiers progress from Moral Sensitivity to Moral Judgment and Moral Agency, scenarios become increasingly complex and cognitively demanding, reflecting a progression through Rest’s moral development model (moral sensitivity, moral judgment, motivation and character) (Narvaez and Rest 1995), Kohlberg’s Stages of Moral Development (pre-conventional, conventional, post-conventional stage) (Kohlberg 1971), and Normative Ethics (Kagan 2018).
Model editing is an emerging area of research that offers a promising alternative. It allows efficient and targeted modifications of language models while minimizing disruptions to their overall knowledge and capabilities (Meng et al. 2022; Wang et al. 2024b; Zhang et al. 2024). While existing work on model editing has focused primarily on updating factual knowledge, its demonstrated effectiveness in making precise and accurate changes to factual knowledge motivates us to extend this idea to ethical behavior steering and introduce the concept of Behavior Editing, which enables directional steering of agent behavior through editing either the agent’s actions or its underlying moral judgments. Behavior Editing not only allows for scenario-specific behavioral adjustments but also enables more extensive shifts in the agent’s global moral alignment. Crucially, this capacity to steer behavior operates in both directions as shown in Figure 1: it can be used to promote benevolent behavior or to induce harmful behavior. In this sense, Behavior Editing is a double-edged sword, simultaneously enabling beneficial interventions and posing significant safety risks.
To systematically investigate this emerging paradigm, we introduce BehaviorBench, a multi-tier benchmark designed to evaluate the effectiveness of behavior editing techniques. Grounded in psychological theories including Moral Foundations Theory (Graham et al. 2013), Stages of Moral Development (Kohlberg 1971), Normative Ethics (Kagan 2018), and Rest’s Four Component Model (Narvaez and Rest 1995), BehaviorBench includes a set of representative model editing approaches, an evaluation framework, and a curated collection of ethical scenarios and dilemmas as summarized in Table 1. Through comprehensive experiments across agents based on both proprietary and open-weight LLMs, we demonstrate that Behavior Editing enables reliable steering of agent behavior across diverse scenarios. Our findings provide new insights into the promise and perils of Behavior Editing, highlighting its potential to enable safer, more ethical agent systems while also revealing the serious risks associated with its misuse.
@@ 2.1 Model Editing
Model editing, also known as knowledge editing (Wang et al. 2024b), has emerged as a critical area of research for modifying LLMs without the need for large datasets or costly retraining. These methods enable precise and efficient updates to specific knowledge while preserving the overall model capabilities. Model editing approaches can be broadly categorized into two groups: parameter-modifying methods and parameter-preserving methods. Parameter-modifying methods alter the model’s internal weights to encode new knowledge. This includes Locate-then-edit techniques such as ROME (Meng et al. 2022), which identifies relevant knowledge within the model before applying targeted modifications, and constrained Fine-Tuning (Zhu et al. 2020; Zhang et al. 2024), which selectively fine-tunes specific layers of the model while minimizing unintended changes. In contrast, parameter-preserving methods such as In-Context Editing (ICE) (Zheng et al. 2023) embed the desired information directly into the input context at inference time, enabling flexible and temporary behavior shifts without altering the underlying model.
These editing techniques have demonstrated effectiveness in updating factual knowledge (Zhang et al. 2024; Huang et al. 2025). However, they also raise safety concerns, particularly the risk of injecting harmful content (Wang et al. 2024a; Chen et al. 2024). As such, model editing presents both opportunities and challenges. Compared to prior work, we demonstrate that model editing can be an effective and precise method for steering LLM-based agents toward specific actions. Furthermore, we find that behavior editing has a substantial impact on the model’s global moral alignment.
@@ 2.2 Ethical Behavior of LLM-based Agents
Datasets relevant to machine ethics include Social Chemistry (Forbes et al. 2020), MoralChoice (Scherrer et al. 2023), ETHICS (Hendrycks et al. 2020), and Jiminy Cricket (Hendrycks et al. 2021). Building on these foundations, our proposed BehaviorBench systematically organizes and enhances existing datasets to evaluate agents across three tiers of moral competence, grounded in psychological and philosophical theory: Moral Sensitivity (recognition of ethical issues), Moral Judgment (reasoned decision-making and justification), and Moral Agency (deliberation and action in ambiguous dilemmas). In comparison to prior benchmarks that primarily emphasize harm avoidance, BehaviorBench offers a more comprehensive assessment of ethical reasoning. It also incorporates a range of normative ethical theories, including deontology, utilitarianism, virtue ethics, theories of justice, and commonsense morality, following the principles of Normative Ethics (Kagan 2018). This multidimensional approach facilitates a more nuanced understanding of how editing techniques can steer agent behavior toward specific ethical orientations.
@@
Various approaches have been proposed to instill ethical constraints and guide the behavior of LLM-based agents. Recent work has focused primarily on alignment techniques. Techniques such as reinforcement learning from human feedback (RLHF) (Ouyang et al. 2022) enable LLMs to align closely with human ethical intuitions by learning from explicit human judgments. Constitutional AI (Sharma et al. 2025; Bai et al. 2022) extends this by allowing models to critique their own outputs against a set of principles. Other methodologies leverage rule-based ethical frameworks or explicit fine-tuning to embed ethical principles directly into model parameters (Choi, Kim, and Lee 2024).
@@
Unlike traditional methods that tune the overall behavior of the model, model editing offers more precise intervention by targeting specific knowledge or response patterns within the parameters of the model while preserving other capabilities (Meng et al. 2022; Zhang et al. 2024). Where RLHF and similar approaches require extensive datasets, computationally expensive retraining, and human involvement, behavior editing can implement targeted behavioral changes with significantly lower computational overhead. This surgical precision makes model editing particularly well-suited for steering ethical behavior in complex scenarios where general alignment techniques might over-constrain agent behavior or fail to address nuanced ethical distinctions.
@@ 3.1 Problem Formulation
The goal of Behavior Editing is to precisely and efficiently steer the behavior of LLM-based agents while preserving their general capabilities. Behavior Editing has two primary directions: Benevolent Behavior Editing, which enhances positive behaviors by steering agents toward more friendly, helpful, and altruistic responses, and Malicious Behavior Editing, which deliberately introduces harmful behaviors to manipulate agents into behaving selfishly, thereby compromising their moral and safety alignment.
@@
Behavior Editing operates on a structure analogous to a knowledge tuple $(s,r,o)$, where traditionally $s$, $r$, and $o$ denote the subject, relation, and object, respectively. However, in the context of Behavior Editing, the interpretations of $s$ and $r$ vary depending on the specific editing settings. We distinguish between two primary categories: Behavior-as-target editing and Judgment-as-target editing, both of which can be represented using the same tuple notation for consistency. In the Behavior-as-target setting, the goal is to modify a behavior exhibited in a given moral scenario. This is formalized as transforming an original tuple $(s,r,o)$, where $s$ is a hypothetical moral scenario, $r$ is the relation to behavior, and $o$ is the behavior under that scenario, into a new tuple $(s,r,o^{*})$ that reflects the edited behavior. Here, the scenario remains constant while the behavior changes. In contrast, Judgment-as-target editing focuses on altering the moral judgment associated with a given behavior. This is represented as transforming the tuple $(s,r,o)$, where $s$ denotes a behavior, $r$ is the relation to moral evaluation, and $o$ is the original moral judgment, into $(s,r,o^{*})$, where $o^{*}$ is the updated judgment. In both cases, an editing operation can be compactly expressed as $e=(s,r,o,o^{*})$, capturing the transformation from the original to the modified output.
@@
To analyze and modify an agent’s behavior in a given scenario, the scenario must first be converted into a natural language question $x$, to which the agent responds with an answer $y$. This input-output pair is associated with a behavior tuple $(s,r,o)$. The input space corresponding to an edit is denoted as $\mathcal{X}_{e}=I(s,r)$, where $I$ maps the scenario and relation to a set of relevant inputs. The original output space is defined as $\mathcal{Y}_{e}=O(s,r,o)$, and the target output space after editing is represented as $\mathcal{Y}_{e}^{*}=O^{*}(s,r,o^{*})$. For a single edit $e$ with input space $\mathcal{X}_{e}$, the objective of Behavior Editing is to transform the original outputs $\mathcal{Y}_{e}$ into the target outputs $\mathcal{Y}_{e}^{*}$. When considering a set of edits $\mathcal{E}=\{e_{1},e_{2},\ldots\}$, the combined input space is $\mathcal{X}_{\mathcal{E}}=\bigcup_{e\in\mathcal{E}}\mathcal{X}_{e}$, and the corresponding original and target output spaces are $\mathcal{Y}_{\mathcal{E}}=\bigcup_{e\in\mathcal{E}}\mathcal{Y}_{e}$ and $\mathcal{Y}_{\mathcal{E}}^{*}=\bigcup_{e\in\mathcal{E}}\mathcal{Y}_{e}^{*}$, respectively.
@@
The overarching goal of Behavior Editing is to modify an LLM-based agent, initially represented as a function $f:\mathcal{X}\rightarrow\mathcal{Y}$, into a new function $f^{*}:\mathcal{X}\rightarrow\mathcal{Y}^{*}$, such that the edited model generates the target behavior for inputs in $\mathcal{X}_{\mathcal{E}}$ while preserving its behavior on all other inputs. The optimization aims to minimize the discrepancy between the edited output $f^{*}(x)$ and the desired behavior $y^{*}$, as measured by a loss function $\mathcal{L}$. At the same time, the editing must maintain consistency across all inputs outside the editing set, ensuring that $f^{*}(x)=f(x)$ for all $x\in\mathcal{X}\setminus\mathcal{X}_{\mathcal{E}}$. This leads to the following constrained optimization objective:
@@
$\text{s.t. }f^{*}(x)=f(x),\quad\forall x\in\mathcal{X}\setminus\mathcal{X}_{\mathcal{E}}$